Hi I need help,
I need to expand and simplify this question:
2(5-x2) +5 (x-1)2

Answer by rothauserc(4718)

2(5-x^2) +5 (x-1)^2 =
:
2(5-x^2) +5(x^2 -2x +1) =
:
10 -2x^2 +5x^2 -10x +5 =
:
*************
3x^2 -10x +15
*************
: